Crane Company purchases equipment on January 1, Year 1, at a cost of $474,000. The asset is expected to have a service life of 12 years and a salvage value of $42,660.

Compute the amount of depreciation for each of Years 1 through 3 using the straight-line depreciation method. (Round answers to 0 decimal places, e.g. 5,125.)

Depreciation for Year 1

$

Depreciation for Year 2

$

Depreciation for Year 3

$

Compute the amount of depreciation for each of Years 1 through 3 using the sum-of-the-years'-digits method.

Depreciation for Year 1

Depreciation for Year 2

Depreciation for Year 3

Compute the amount of depreciation for each of Years 1 through 3 using the double-declining-balance method. (Round depreciation rate to 2 decimal places, e.g. 15.84%. Round answers to 0 decimal places, e.g. 45,892.)

Depreciation for Year 1

Depreciation for Year 2

Depreciation for Year 3
Answer #1

Straight line method = (474,000-42,660)/12 = 35945

Depreciation for year 1 35945
Depreciation for year 2 35945
Depreciation for year 3 35945

SYD = 12*(12+1)/2 = 78

Depreciation for year 1 (12/78)*(474,000-42,660) 66360
Depreciation for year 2 (11/78)*(474,000-42,660) 60830
Depreciation for year 3 (10/78)*(474,000-42,660) 55300

Double declining rate = 200/Useful life = 200/12 = 16.67%

Depreciation for year 1 (474,000*16.67%) 79015
Depreciation for year 2 (474,000-79015)*16.67% 65,843
Depreciation for year 3 54867